# Polarization-encoded quantum key distribution with a room-temperature telecom single-photon emitter

### v1

@@ I Introduction
Many materials have been investigated as solid-state SPSs. Examples include semiconductor quantum dots [11, 12, 13, 14] and nitrogen-vacancy color centers [15, 16]. Unfortunately, the emission wavelengths of most SPSs are shorter than the telecom band [17, 18, 19], where dedicated fibers [20] or additional frequency conversion steps [21, 22] are required. Additionally, many of these sources require cryogenic operation [23, 24]. In particular, reliable telecom-band SPSs are desirable due to their inherent suitability with the existing deployed fiber links [17, 25, 26].
@@ II.3 Key Rates
Fig. 5 (a) records the variation in QBER and sifted key rate for signals transmitted over the deployed fiber, with an integration time of 20 s. The H and V polarizations were pre-aligned to the PSPs of the channel as we established earlier, which resulted in a HV QBER of 1.7% and DA QBER of 8.3%. A raw key rate of 1349.6 bps with an average QBER of 5.0% was obtained over 7 hours without active feedback. With the experimental parameters of the devices and the assumed variables summarized in Table. 2, a finite-key analysis could be carried out to bound the amount of secure key to be $\epsilon_{sec}$-secret and $\epsilon_{cor}$-correct [35]. In this experiment, the HV and DA bases had the same preparation and measurement probability, where a secure key rate of 247.3 bps was observed with finite key analysis. However, considering the nature of unbalanced QBER caused by PMD, an unbalanced ratio of bases could be designed to maximize the key rate. A secure key rate of 585.9 bps could be achieved under $p_{z}=0.997$ and $p_{x}=0.003$, which well approaches the asymptotic limitation from GLLP [36, 37] considering our channel and device characteristics as shown in Fig. 5(b). The result for the 32.5 km spool is also indicated with a reasonably low QBER of 3.2% and a secure key rate of 50.4 bps.

### v2

@@ I Introduction
Many materials have been investigated as solid-state SPSs. Examples include semiconductor quantum dots [11, 12, 13, 14] and nitrogen-vacancy color centers [15, 16]. Reliable telecom-band SPSs are particularly desirable for QKD applications due to their compatibility with existing deployed fiber links. However, most SPSs emit at wavelengths shorter than the telecom band [17, 18, 19], where dedicated fibers [20] or additional frequency conversion steps [21, 22] are required. Moreover, many telecom-band SPSs require cryogenic cooling systems [23, 24, 25], limiting their commercial feasibility for QKD. A GaN-based SPS operating at room temperature within the telecom band[26] is therefore an excellent quantum source for metropolitan QKD [27]. Alternatively, SPSs outside the telecom band, such as hBN, which also operate at room temperature, could enable QKD over free-space links [28], although they are not compatible with fiber-based communication.
@@ II.3 Key Rates
Fig. 5 (a) records the variation in QBER and sifted key rate for signals transmitted over the deployed fiber, with an integration time of 20 s. The D and A polarizations were pre-aligned to the PSPs of the channel as we established earlier, which resulted in a DA QBER of 1.7% and LR QBER of 8.3%. A raw key rate of 1349.6 bps with an average QBER of 5.0% was obtained over 7 hours without active feedback. With the experimental parameters of the devices and the assumed variables summarized in Table. 2, a finite-key analysis could be carried out to bound the amount of secure key to be $\epsilon_{sec}$-secret and $\epsilon_{cor}$-correct [37]. In this experiment, the DA and LR bases had the same preparation and measurement probability, where a secure key rate of 247.3 bps was observed with finite key analysis. However, considering the nature of unbalanced QBER caused by PMD, an unbalanced ratio of bases could be designed to maximize the key rate. A secure key rate of 585.9 bps could be achieved under $p_{z}=0.997$ and $p_{x}=0.003$, which well approaches the asymptotic limitation from GLLP [38, 39] considering our channel and device characteristics as shown in Fig. 5(b). The result for the 32.5 km spool is also indicated with a reasonably low QBER of 3.2% and a secure key rate of 50.4 bps.
@@ Acknowledgements.
Author Contributions
X.Z. and H.Z. performed the experiment and coding and analyzed the data. J. E. and M. M. implemented the methods on the extraction and measurement of the SPS. J. A. G. and C. R. M. developed ideas for the PMD section. W. G. and A. L. supervised all work, as well as contributed to deciding the experiment scheme and setting up the physical models. All authors discussed the results and contributed to the manuscript. Another experiment was conducted parallelly using a complementary method, verifying the feasibility of time-bin encoding protocols with the GaN SPS.